A particle moves along x axis is such way that its coordinate x varies with time T according to expression x=2-5t+6t^2(square) the initial velocity of according of particlea.-3,-6b.-3,12c.-5,12d.-5,6

Grade:11

1 Answers

Arun
6 years ago
Particle is moving along x-axis so displacement is also along x-axis 

Now , ⇒ x = 6t²-5t+2

since , V = dx/dt

Differentiating w.r.t. 't'

⇒ v = 6 × 2t - 5

⇒ v = 12t - 5
Hence initial velocity
v = 0-5 = -5 m/sec